A light-emitting diode (LED) connected to a 3.0 V power supply emits 440 nm blue light. The current in the LED is 11 mA , and the LED is 71 % efficient at converting electric power input into light power output.

A) How many photons per second does the LED emit?
Answer #1

Energy, E = hc/lambda

E = (6.63 x 10^-34 x 3 x 10^8)/(440 x 10^-9) = 4.52 x 10^-19 J

Electric energy = VI = 3 x 0.011 = 0.033 W

Converted energy = 0.033 x 0.71 = 0.023 J

No. of photons, n = 0.023/(4.52 x 10^-19) = 5.18 x 10^16 photons

Comment in case any doubt please rate my answer....